1. The jewelry department had net sales of $10,000. Operating expenses were 43% of the sales. Cost of goods sold was $550. What is the profit/loss for the department?
Answer #1

Girss margin = Net sales - Cost of goods sold = 10,000 - 550 = 9,450

Expenses = 43% of 10,000 = 4,300

Profit/Loss = Gross margin - Expenses = 9450 - 4300 = $ 5150

Do give a positive rating if you think this answer helped